REU: SCHEDULE FOR WEEKS SEVEN AND EIGHT

Date  SESAME May (cats) Kiselev (QM) Brock (geo)
Monday, July 29  9–12, 1–4 4:15 – 5:45
Tuesday, July 30  9–12, 1–4 12:10 – 12:50 4:15 – 5:45
Wednesday, July 31 9–12, 1–4 12:10 – 12:50  4:15 – 5:45
Thursday, August 1 9–12, 1–4 12:10 – 12:50 4:15 – 5:45
Friday, August 2  9–12, 1–4 12:10 – 12:50  4:15 – 5:45
Monday, August 5  9–12, 1–4 12:10 – 12:50  4:15 – 5:45
Tuesday, August 6  9–12, 1–4 12:10 – 12:50 4:15 – 5:45
Wednesday, August 7 9–12, 1–4 12:10 – 12:50  4:15 – 5:45
Thursday, August 8 9–12, 1–4 12:10 – 12:50 4:15 – 5:45
Friday, August 9  9–12, 1–4 12:10 – 12:50

I urge graduate student counsellors to offer daily afternoon problem sessions, between 1:00 and 4:00, for the YSP participants (and SESAME participants who work in the mornings) on any and all of the courses to date. All undergraduate participants should feel free to e-mail graduate students to set up one-on-one or small group meetings, and I remind undergraduates that we have many graduate student volunteers willing to mentor directed reading programs between the end of the REU and the beginning of the fall quarter.

Here is the schedule in summary form:

Seventh week:

SESAME: 9–12, 1–4 M – F May 4:15 – 5:45 Monday, 12:10–12:50 Tuesday – Friday Kiselev 4:15 – 5:45 Tuesday and Thursday Brock 4:15 – 5:45 Wednesday and Friday

Eighth week:

SESAME: 9–12, 1–4 M – F May 12:10–12:50 Monday – Friday Kiselev 4:15 – 5:45 Tuesday and Thursday Brock 4:15 – 5:45 Monday and Wednesday Party?! 5:00 – ?:?? Friday